# Josephson tunnel junction arrays and Andreev weak links: linked by a single energy-phase relation

### v1

@@ III What insights can we draw from these correspondences?
Cooper-pair-parity Protected Qubits: The same kind of interferometers as in the diodes above have been considered for qubits that are protected by conservation of Cooper pair parity, both in the SISIS case (known as the rhombus) [31, 32] and in SNS junctions [33, 34, 35]. When the two interferometer arms are identical and the flux is tuned to half a flux quantum, the odd-harmonic terms in the EPR cancel out, $\cos(n_{\mathrm{odd}}\varphi)+\cos(n_{\mathrm{odd}}(\varphi+\pi))=0$, while the even-harmonic terms remain, $\cos(n_{\mathrm{even}}\varphi)+\cos(n_{\mathrm{even}}(\varphi+\pi))=2\cos(n_{\mathrm{even}}\varphi)$. Because the EPR harmonics are related to Cooper pair transfers, the interferometer now only mediates transfers of even numbers of Cooper pairs between the reservoirs. Therefore, the parity of the number of transferred Cooper pairs becomes a conserved quantity, which may be useful for designing a qubit with long coherence times.
@@ IV Discussion
Intermediate coupling: The cases considered in Sec. II, while instructive and useful, are the theoretical limiting cases. At intermediate coupling, the EPRs no longer strictly follow Eq. (1). For SNS, one must account for the influence the above-gap continuum in the superconducting density of states, which itself contributes to the supercurrent and distorts the dispersion of bound state [56, 7, 57]. For SISIS, one must consider more carefully the dynamics of the charge on the island, such as by including additional charge states [10]. Also, we are not aware of direct mathematical correspondences when the number of series tunnel junctions is more than two [58] or the SNS junction is not short [59].
@@
Gate tunability: Finally, SNS implementations with low electron density normal regions, such as semiconductors and 2d materials, afford in situ tunability of the junction strength $E_{0}$ and effective transparency $\tau_{\mathrm{eff}}$ with an electrostatic gate voltage [71, 72, 73, 74, 75], as depicted in Fig. 2(b). Such in situ tunability is a powerful tool for validating theoretical ideas, uncovering experimental realities, and parametric control, as ha been taken advantage of in the past through magnetic flux tunability of loops containing SIS junctions [76, 77]. The charge tunability made possible by SNS junctions adds a complementary control axis with low cross-talk to flux-tuned circuit elements and lower power dissipation elsewhere in the cryogenic setup. In a recent experimental work [78], it has been demonstrated that the energy-phase relation of two voltage-controlled SNS junctions in series follow Eq. (1), supporting a model of two conventional JJs in series [17]. Example demonstrations of practical applications include modulation between optimal operation points [79, 80, 81, 82, 83] and activation of dynamic effects like amplification [84, 85, 86, 87]. Charge control over the junction inductance also may enable novel circuit couplings for nonreciprocal devices such as the gyrator and derivative qubit concepts [88]. In principle, gate-tunability is also possible for SISIS junctions in the weak and intermediate coupling regimes [10]; however, challenges include achieving useful gate-modulation amplitudes and reliable device handling. We also remark that protected qubit concepts, like the rhombus, often require high circuit symmetry, placing stringent requirements on Josephson junction fabrication accuracy, which is only reliable at the % level in the state of the art [89, 90, 91]. Gate tunability can help resolve that issue for individual devices, at the cost of introducing a new noise channel.

### v2

@@ III What insights can we draw from these correspondences?
Cooper-pair-parity Protected Qubits: The same kind of interferometers as in the diodes above have been considered for qubits that are protected by conservation of Cooper pair parity, both in the SISIS case (known as the rhombus) [31, 32] and in SNS junctions [33, 34, 35] (and also recently in a hybrid structure [36]). When the two interferometer arms are identical and the flux is tuned to half a flux quantum, the odd-harmonic terms in the EPR cancel out, $\cos(n_{\mathrm{odd}}\varphi)+\cos(n_{\mathrm{odd}}(\varphi+\pi))=0$, while the even-harmonic terms remain, $\cos(n_{\mathrm{even}}\varphi)+\cos(n_{\mathrm{even}}(\varphi+\pi))=2\cos(n_{\mathrm{even}}\varphi)$. Because the EPR harmonics are related to Cooper pair transfers, the interferometer now only mediates transfers of even numbers of Cooper pairs between the reservoirs. Therefore, the parity of the number of transferred Cooper pairs becomes a conserved quantity, which may be useful for designing a qubit with long coherence times.
@@ IV Discussion
Intermediate coupling: The cases considered in Sec. II, while instructive and useful, are the theoretical limiting cases. At intermediate coupling, the EPRs no longer strictly follow Eq. (1). For SNS, one must account for the influence the above-gap continuum in the superconducting density of states, which itself contributes to the supercurrent and distorts the dispersion of bound state [57, 7, 9, 58]. Additionally, violation of the Andreev approximation (e.g., if the bands disperse significantly within the energy window of $2\Delta$) results in deviations as well [59]. For SISIS, one must consider more carefully the dynamics of the charge on the island, such as by including additional charge states [10]. Also, we are not aware of direct mathematical correspondences when the number of series tunnel junctions is more than two [60] or the SNS junction is not short [61].
@@
Gate tunability: Finally, SNS implementations with low electron density normal regions, such as semiconductors and 2d materials, afford in situ tunability of the junction strength $E_{0}$ and effective transparency $\tau_{\mathrm{eff}}$ with an electrostatic gate voltage [73, 74, 75, 76, 77], as depicted in Fig. 2(b). Such in situ tunability is a powerful tool for validating theoretical ideas, uncovering experimental realities, and parametric control, as ha been taken advantage of in the past through magnetic flux tunability of loops containing SIS junctions [78, 79]. The charge tunability made possible by SNS junctions adds a complementary control axis with low cross-talk to flux-tuned circuit elements and lower power dissipation elsewhere in the cryogenic setup. In recent experimental works [80, 36], it has been demonstrated that the energy-phase relation of two voltage-controlled SNS junctions in series follow Eq. (1), supporting a model of two conventional JJs in series [17]. Example demonstrations of practical applications include modulation between optimal operation points [81, 82, 83, 84, 85] and activation of dynamic effects like amplification [86, 87, 88, 89]. Charge control over the junction inductance also may enable novel circuit couplings for nonreciprocal devices such as the gyrator and derivative qubit concepts [90]. In principle, gate-tunability is also possible for SISIS junctions in the weak and intermediate coupling regimes [10]; however, challenges include achieving useful gate-modulation amplitudes and reliable device handling. We also remark that protected qubit concepts, like the rhombus, often require high circuit symmetry, placing stringent requirements on Josephson junction fabrication accuracy, which is only reliable at the % level in the state of the art [91, 92, 93]. Gate tunability can help resolve that issue for individual devices, at the cost of introducing a new noise channel.